If a farmer can see 12 miles from a firetower, How many square miles can he see?
Is 75.36 right?

Answer by checkley71(8403)

WHAT YOU HAVE IS A CIRCLE THAT HAS A RADIUS=12 MILES.
AREA=PI*R^2
AREA=3.14*12^2
AREA=3.14*144
AREA=452.16 SQUARE MILES IS THE ANSWER.